Suppose that the coordinates of R be (3,4,12), what are the direction cosines of
(
O
R
)
(OR) ƒ—?
1
Expert's answer
2020-06-09T18:16:48-0400

"\overrightarrow{OR}=(3-0)\hat{i}+(4-0)\hat{j}+(12-0)\hat{k}=" "3\hat{i}+4\hat{j}+12\hat{k}"

"|\overrightarrow{OR}|=\sqrt{3^2+4^2+12^2}=13"

Unit vector "=\frac{3}{13}\hat{i}+\frac{4}{13}\hat{j}+\frac{12}{13}\hat{k}"

Direction cosines "\equiv (\frac{3}{13},\frac{4}{13},\frac{12}{13})"